# Toward Low-Latency, High-Fidelity Calibration of the LIGO Detectors with Enhanced Monitoring Tools

### v1

@@ 3 Anticipated Challenges for LIGO Calibration in O5
The past 10 years has brought significant advancements in LIGO calibration efforts and has moved us from a place of providing final calibration data products (strain data and accompanying systematic error estimates) on timescales of months towards timescales of hours. Although, the process of reviewing and certifying calibration products still takes weeks or months. Moving forward, the LIGO calibration effort will need to continue to evolve in tandem with ongoing improvements to the LIGO detectors. Detector upgrades will impact the DARM loop that is at the heart of calibration, and requirements on the accuracy and latency of the calibrated data will continue to tighten as louder and more gravitational-wave sources become observable.
The planned detector upgrades for O5 that will significantly impact calibration include the implementation of a Balanced Homodyne Detection (BHD) readout scheme [48], replacement of the test masses and an increase to the circulating arm power from 400 kW to 750 kW [49]. The BHD readout scheme will introduce an adjustable homodyne angle, the audio-band demodulation quadrature of the GW readout, which will need to be included in the sensing function model [14]. The higher circulating arm power will lead to larger thermal transients in the sensing response of the detectors each time the detectors transition from an unlocked and uncontrolled (cold) state to the full-power (hot) operating state. The replacement of test masses will require an update to the models of suspension dynamics, which directly impacts the actuation function. The test mass upgrade also has the potential for resetting the number and location of point absorbers, which influence both the sensing response and, indirectly, the amount of cross-coupling in the actuation dynamics.
The planned upgrade to circulating arm power of 750 kW is likely to be particuarly impactful on the calibration in O5. As the arm cavity mirrors heat up during the thermalization period, they experience surface deformation due to absorption from the main interferometer beam. This deformation will affect the frequency-dependent sensing response, $C(f)$. However, the current sensing function model is not sophisticated enough to account for thermalization effects in real time. The detector strain reconstructed during each thermalization period will then contain errors due to the mismatch between the sensing function model used for reconstruction and the true interferometer sensing function during the thermalization period. The interferometer is said to be thermalized once the the circulating arm power reaches roughly 95% of its nominal operating value and the thermal transient is observed to resign. In O4, that period was roughly 1.5-2 hours. Figs. 3 and 4 in section 4.2 show the effect of thermalization on the systematic error of the calibrated data during O4.
@@
Another approach to addressing the negative impacts of thermalization on the sensing function is to reduce the physical impact of thermalization on the detector response. There are ongoing efforts to deploy additional heating actuators that will effectively heat the test mass surfaces to mitigate the effects of thermalization during power up. These are the The FROnt Surface Type Irradiator (FROSTI) [50] and Central Heater for Transient Attenuation (CHETA) [8]. In particular, the CHETA actuator promises to significantly mitigate the majority of calibration-affecting thermalization issues by using a CO2 laser source pointed at the cavity-facing surfaces of the input and end test masses to maintain the mirrors in a “thermalized” state via absorption at times when the interferometer is not locked (i.e. when the circulating arm power is $\sim$0 W). Once successfully deployed, the CHETA project will significantly shrink the gap between the “cold” and “hot” interferometer states, rendering thermalization effects negligible.
As discussed in section 2.4, estimates for LIGO calibration systematic error rely on measurements of the DARM loop taken at specific moments in time. For gravitational wave signals that are broadband and transient in nature, e.g., binary neutron star coalesence, the broadband estimate of calibration systematic error can be mapped in a straightforward way to this waveform, and it is safe to assume the calibration systematic error is static during the measurement of the gravitational wave signal. On the other hand, for signals that are narrowband and long-duration in nature, e.g., rapidly-rotating neutron stars emitting a weak, continuous, quasi-monochromatic gravitational wave signal, the calibration systematic error might not be well-accounted by the broadband, static estimate. Continued work is required in order to properly estimate calibration systematic error for narrowband signals and to integrate those time-varying estimates into posterior distributions of source parameter estimates.
@@ 4 Advancements in LIGO Calibration Accuracy, Latency, and Monitoring
Due to the increased rate of GW detections [7], the improvements in pipeline reliability discussed below, the development of a real-time monitoring system on the low-latency calibrated data, and the development of a low-latency estimate of calibration error and uncertainty, the decision was made at the beginning of O4 to use the low-latency strain data as the finalized calibrated strain data product whenever possible. Prior to O4, final published results of analyses were based on a second, high-latency calibration produced from archived data, months after initial data acquisition. This was done primarily to ensure accurate knowledge of the calibration models and a well-informed estimate of the systematic error and uncertainty in those models. Since the beginning of O4, a high-latency calibration is produced only occasionally for time periods with known significant errors.
@@ 4.1 Low-latency Calibration Pipeline Improvements
The current low-latency calibration pipeline is divided into two stages, the first of which is accomplished using the front-end computers used to control the detector. The front-end calibration uses infinite impulse response (IIR) filtering to compute an estimate of the strain in each detector. The front-end calibration pipeline’s exclusive use of IIR filters and limitation to timestamping data only in real time lead to deficiencies in the pipeline’s ability to accurately represent the detector response. The second stage of calibration is accomplished in the gstlal calibration pipeline [44, 51], a GStreamer-based pipeline that wraps LIGO Algorithm Library (LAL[52] software with GStreamer’s audio streaming capability. The gstlal calibration pipeline is currently used to produce calibrated data in real time using partially calibrated data from the front end as input, as well as in high latency using archived raw data as input. The gstlal calibration pipeline uses time-domain finite impulse response (FIR) filters and is therefore able to apply calibration models with arbitrary frequency dependence and produce strain data that is independent of the pipeline’s start time, a necessity for exactly reproducing strain data. The gstlal calibration pipeline also computes a bitwise state vector whose primary purpose is to indicate times when the strain data is suitable for analysis.
@@ 4.1.1 Improvements to Calibration Pipeline Latency
The contribution of the calibration pipeline to the latency of LIGO strain data has been monitored since the beginning of O2, when the calibration latency was $\sim$13 seconds. Due primarily to two improvements, the calibration latency was reduced to $\sim$3 seconds by the beginning of O3 and has remained at this values throughout O4. One source of latency that was eliminated was caused by anti-aliasing and anti-imaging filters associated with downsampling and upsampling processes in the measurement, calculation, and application of the time-varying parameters of the DARM model. Each of the two such resampling processes added $\sim$6 seconds of latency. This $\sim$12-second latency was eliminated by including an option in the resampling algorithm to simply shift the timestamps of the resampled data. With no latency contribution from this branch of the pipeline, the FIR filters, which occupy a parallel branch of the pipeline, became the dominant source of pipeline latency.
The remaining latency has been minimized by significant improvements to the quality of the FIR filters used in the gstlal calibration pipeline to apply the models $1/C^{\rm(model)}$ and $A^{\rm(model)}$. The FIR filters are computed by taking the inverse Fourier transform of a model that is represented in the frequency domain. The resulting time-domain filter is then windowed to optimize its frequency response. Due to seismic noise at low frequencies, the filters include a very aggressive high-pass rolloff below 10 Hz. Achieving adequate high-pass filtering without incurring a large latency in the strain data is one of the primary challenges in FIR filter design. During O2, the filters were windowed using a Tukey window with a cosine fraction of 0.9. Using this window, the high-pass filtering stipulation required long filters that incurred 4 seconds of latency. By the beginning of O3, a more fine-tuned window design had reduced the filter latency to 2 seconds while also improving the quality of the high-pass filter. Currently, a Slepian window is used, resulting in filters with better attenuation of low-frequency noise and tunable frequency resolution. Small delays associated with data transfer into and out of the pipeline and with all the computations done in the pipeline add $\sim$1 second, leading to the current total latency of $\sim$3 seconds.
@@ 4.1.2 Improvements to Pipeline Reliability and Robustness
Since O1 and O2, the calibration pipeline has also become significantly more reliable and robust. One of the main issues that has been thoroughly addressed was caused by frequent dropouts of raw data at the input to the gstlal calibration pipeline. These dropouts can occur in any subset of the $\sim$100 input channels and can vary in length from seconds to hours, but they rarely exceed a few seconds during observation. A method was introduced at the beginning of O2 to fill data dropouts in with zeros. The calibration state vector indicates times when these dropouts corrupt the strain data. By the beginning of O4, more flexible options have been included in order to minimize loss of data quality in the output. Dropouts can be filled with: (1) zeros - the most common option, indicating data that should not be used in strain production if possible, or that the strain data computed is not suitable for analysis, (2) user-defined nonzero values - useful for instances in which the pipeline needs to handle data dropouts in a different manner that actual zeros in a channel, or (3) the last good value - useful for channels containing constants that rarely change. Additionally, during real-time calibration, the pipeline uses a timer to determine when a dropout is occurring. Prior to this improvement, made during O3, the pipeline could only identify a dropout based on the timestamp of the first input data after the dropout ended. Long dropouts therefore caused the pipeline output to fall far behind real time, causing the delayed strain data to be dropped in the data aggregation processes that follow calibration. This is no longer an issue, as the pipeline continuously fills in missing data during long dropouts.
@@ 4.1.3 Development of an Exact Solution for Time Dependence in the Calibration Model
The methods currently used to compute time-dependent parameters in $C^{\rm(model)}$ and $A^{\rm(model)}$ rely on simplifying approximations, resulting in $\sim$1% errors in the estimates of time-dependent parameters. The out-of-phase addition of the actuation and sensing components of $h_{\rm det}^{\rm(model)}$ shown in (6) inflates these errors in the strain data at some frequencies. An “exact” solution, described in chapter 5 of [54] has been developed in the gstlal calibration pipeline but is not yet implemented in the production of strain data as review and validation is still ongoing. This method will likely be implemented in O5, eliminating a systematic error that is currently uncompensated in the production of strain data. These systematic errors are largely accounted for—though not explicitly modeled—when estimating the overall calibration error and uncertainty budget in O4. The implementation of the exact solution may also produce more consistent improvements to calibration accuracy with the application of the full time-dependent model.
@@ 4.1.4 Subtraction of Spectral Lines
Searches for GWs are negatively impacted by narrow-band noise sources in the strain data [55]. Some narrow-band noise sources in the form of persistent, sinusoidal signals, known as calibration lines, are intentionally added via the Pcal lasers and the actuation system, in part to the measure the time-dependent parameters of the calibration models during observation. A method for subtracting spectral lines, including the calibration lines, in low latency was implemented in the gstlal calibration pipeline starting at the beginning of O3. The method relies on the use of witness channels, which are correlated to the signal to be subtracted but insensitive to GWs. This is intended to prevent the unintentional subtraction of GWs. The method is carried out by first demodulating strain $h(t)$ and the witness channel $w_{j}(t)$ at the frequency of the $i^{\rm th}$ line $f_{i}$ to find complex values $\tilde{h}(f_{i})$ and $\tilde{w}_{j}(f_{i})$ which encode the amplitude and phase of $h$ and $w_{j}$ at that frequency. A transfer function value is then computed at $f_{i}$ using a running median as
@@
During the review of calibrated strain data, several issues and shortfalls of the low-latency line subtraction performed in the low-latency calibration pipeline were identified. These issues were addressed and fixed before the final release of the strain data. One problem that was addressed during O4 is a narrow-band systematic error, often referred to as “oversubtraction,” introduced near the line frequencies. This error is caused by the use of a finite amount of data to inform the transfer function values $T_{j}$, leading to an imperfect ability to distinguish between signal that is present in the witness $w_{j}$ and signal that is not. The bandwidth of this error is inversely proportional to the length of the running median, which had been set to 128 seconds during O3. Increasing the length of the running median mitigates the error, but it also hinders the ability to subtract spectral lines when the $T_{j}$ are evolving on shorter timescales. During O4, a method was implemented that allows the subtraction algorithm to dynamically change the length of the running median based on whether significant secular change is detected in the $T_{j}$ before taking the median. The length of the running median is now adjusted between a minimum of 128 seconds and a maximum of 4096 seconds. Fig. 1 demonstrates the subtraction of the 17.1 Hz Pcal line using the 128 second running median, the 4096 second running median, and the adaptive switching method.
Numerous types of occasional line subtraction failures have been caused by data dropouts and by insufficient communication of detector state information to the subtraction algorithm. Many of these have been addressed by more sophisticated handling of data dropouts and by more thoroughly utilizing all available state information. The current method uses multiple metrics to determine when the input data is suitable for measurement of the $T_{j}$ and when the calibration lines are turned on or off. With the addition of the Pcal lines used for calibration monitoring in O4 (discussed below), the line subtraction performed in the low-latency calibration procedure has become increasingly essential to the production of high-quality strain data for downstream analyses. The improvements to the quality and robustness of the line subtraction algorithm have been essential to this achievement.

### v2

@@ 3 Anticipated Challenges for LIGO Calibration in O5
The decade following the discovery of GW150914 has brought significant advancements in LIGO calibration efforts and has moved us from a place of providing final calibration data products (strain data and accompanying systematic error estimates) on timescales of months towards timescales of hours. Although, the process of reviewing and certifying calibration products still takes weeks to months. Moving forward, the LIGO calibration effort will need to continue to evolve in tandem with ongoing improvements to the LIGO detectors. Detector upgrades will impact the DARM loop that is at the heart of calibration, and requirements on the accuracy and latency of the calibrated data will continue to tighten as louder and more gravitational-wave sources become observable.
The planned detector upgrades for O5 that will significantly impact calibration include the implementation of a Balanced Homodyne Detection (BHD) readout scheme [52], replacement of the test masses and an increase to the circulating arm power from 400 kW to 750 kW [53]. The BHD readout scheme will introduce an adjustable homodyne angle, the audio-band demodulation quadrature of the GW readout, which will need to be included in the sensing function model [14]. The higher circulating arm power, if implemented without additional thermal actuation, will lead to larger thermal transients in the sensing response of the detectors each time the detectors transition from a not-operational (cold) state to the full-power (hot) operating state. Thermal actuation systems, described more below, or a more sophisticated sensing model will need to be developed to mitigate these large thermal transients. The replacement of test masses will require an update to the models of suspension dynamics, which directly impacts the actuation function. The test mass upgrade also has the potential for resetting the number and location of point absorbers, which influence both the sensing response and, indirectly, the amount of cross-coupling in the actuation dynamics.
The planned upgrade to circulating arm power of 750 kW is likely to be particularly impactful on the calibration in O5. As the arm cavity mirrors heat up during the thermalization period, they experience surface deformation due to absorption from the main interferometer beam. This deformation will affect the frequency-dependent sensing response, $C(f)$. However, the current sensing function model is not sophisticated enough to account for thermalization effects in real time. The detector strain reconstructed during each thermalization period will then contain errors due to the mismatch between the sensing function model used for reconstruction and the true interferometer sensing function during the thermalization period. The interferometer is said to be thermalized once the the circulating arm power reaches roughly 95% of its nominal operating value and the thermal transient is observed to resign. In O4, that period was roughly 1.5-2 hours. Figs. 4 and 5 in section 4.2 show the effect of thermalization on the systematic error of the calibrated data during O4.
@@
Another approach to addressing the negative impacts of thermalization on the sensing function is to reduce the physical impact of thermalization on the detector response. There are ongoing efforts to deploy additional heating actuators that will effectively heat the test mass surfaces to mitigate the effects of thermalization during power up. These are the The FROnt Surface Type Irradiator (FROSTI) [54] and Central Heater for Transient Attenuation (CHETA) [8]. In particular, the CHETA actuator promises to significantly mitigate the majority of calibration-affecting thermalization issues by using a CO2 laser source pointed at the cavity-facing surfaces of the input and end test masses to maintain the mirrors in a “thermalized” state via absorption at times when the circulating arm power is $\sim$0 W. Once successfully deployed, the CHETA project will significantly shrink the gap between the “cold” and “hot” interferometer states, rendering thermalization effects negligible.
As discussed in section 2.4, estimates for LIGO calibration systematic error rely on measurements of the DARM loop taken at specific moments in time. For gravitational wave signals that are broadband and transient in nature, e.g., a binary neutron star coalescence, the broadband estimate of calibration systematic error can be mapped in a straightforward way to this waveform, and it is safe to assume the calibration systematic error is static during the measurement of the gravitational wave signal. On the other hand, for signals that are narrow-band and long-duration in nature, e.g., rapidly-rotating neutron stars emitting a weak, continuous, quasi-monochromatic gravitational-wave signal, the calibration systematic error might not be well-accounted by the broadband, static estimate. Continued work is required in order to properly estimate calibration systematic error for narrow-band signals and to integrate those time-varying estimates into posterior distributions of source parameter estimates.
@@ 4 Advancements in LIGO Calibration Accuracy, Latency, and Monitoring
Due to the increased rate of GW detections [7], the improvements in pipeline reliability discussed below, the development of a real-time monitoring system on the low-latency calibrated data, and the development of a low-latency estimate of calibration error and uncertainty, the decision was made at the beginning of O4 to use the low-latency strain data as the final calibrated strain data product whenever possible. Prior to O4, final published results of analyses were based on a second, high-latency calibration produced from archived data, months after initial data acquisition. This was done primarily to ensure accurate knowledge of the calibration models and a well-informed estimate of the systematic error and uncertainty in those models. Since the beginning of O4, a high-latency calibration is produced only occasionally for time periods with known significant errors. For O4a and O4b, which spanned approximately 20 months of observing time, a total of about 2 months of data required high-latency recalibration.
@@ 4.1 Low-latency Calibration Pipeline Improvements
The low-latency calibration pipeline used from O1–O4 is divided into two stages, the first of which is accomplished using the front-end computers used to control the detector. The front-end calibration uses infinite impulse response (IIR) filtering to compute an estimate of the strain in each detector. The front-end calibration pipeline’s exclusive use of IIR filters and limitation to timestamping data only in real time lead to deficiencies in the pipeline’s ability to accurately represent the detector response. The second stage of calibration is accomplished in the gstlal calibration pipeline [47, 55], a GStreamer-based pipeline that wraps LIGO Algorithm Library (LAL[56] software with GStreamer’s audio streaming capability. The gstlal calibration pipeline was used from O1–O4 to produce calibrated data in real time using partially calibrated data from the front end as input, as well as in high latency using archived raw data as input. The gstlal calibration pipeline uses time-domain finite impulse response (FIR) filters and is therefore able to apply calibration models with arbitrary frequency dependence and produce strain data that is independent of the pipeline’s start time, a necessity for exactly reproducing strain data. The gstlal calibration pipeline also computes a bitwise state vector whose primary purpose is to indicate times when the strain data is suitable for analysis.
@@ 4.1.1 Improvements to Calibration Pipeline Latency
The contribution of the calibration pipeline to the latency of LIGO strain data has been monitored since the beginning of O2, when the calibration latency was $\sim$13 seconds. Due primarily to two improvements, the calibration latency was reduced to $\sim$3 seconds by the beginning of O3 and has remained at this values throughout O4. To put this latency in perspective, the first preliminary alert for public gravitational-wave events throughout O4 has an expected median latency of $\sim 30$ seconds [57]. One source of latency in the calibration pipeline that was eliminated was caused by anti-aliasing and anti-imaging filters associated with downsampling and upsampling processes in the measurement, calculation, and application of the time-varying parameters of the DARM model. Each of the two such resampling processes added $\sim$6 seconds of latency. This $\sim$12-second latency was eliminated by including an option in the resampling algorithm to simply shift the timestamps of the resampled data. With no latency contribution from this branch of the pipeline, the FIR filters, which occupy a parallel branch of the pipeline, became the dominant source of pipeline latency.
The remaining latency has been minimized by significant improvements to the quality of the FIR filters used in the gstlal calibration pipeline to apply the models $1/C^{\rm(model)}$ and $A^{\rm(model)}$. The FIR filters are computed by taking the inverse Fourier transform of a model that is represented in the frequency domain. The resulting time-domain filter is then windowed to optimize its frequency response. Due to seismic noise at low frequencies, the filters include a very aggressive high-pass roll-off below 10 Hz. Achieving adequate high-pass filtering without incurring a large latency in the strain data is one of the primary challenges in FIR filter design. During O2, the filters were windowed using a Tukey window with a cosine fraction of 0.9. Using this window, the high-pass filtering stipulation required long filters that incurred 4 seconds of latency. By the beginning of O3, a more fine-tuned window design had reduced the filter latency to 2 seconds while also improving the quality of the high-pass filter. In O4, a Slepian window is used, resulting in filters with better attenuation of low-frequency noise and tunable frequency resolution. Small delays associated with data transfer into and out of the pipeline and with all the computations done in the pipeline add $\sim$1 second, leading to the current total latency of $\sim$3 seconds.
@@ 4.1.2 Improvements to Pipeline Reliability and Robustness
Since O1 and O2, the calibration pipeline has also become significantly more reliable and robust. One of the main issues that has been thoroughly addressed was caused by frequent dropouts of raw data at the input to the gstlal calibration pipeline. These dropouts can occur in any subset of the $\sim$100 input channels and can vary in length from seconds to hours, but rarely exceed a few seconds during observation. A method was introduced at the beginning of O2 to fill in the data dropouts with zeros. The calibration state vector indicates times when these dropouts corrupt the strain data. By the beginning of O4, more flexible options were included in order to minimize loss of data quality in the output. Dropouts can be filled with: (1) zeros - the most common option, indicating data that should not be used in strain production if possible, or that the strain data computed is not suitable for analysis, (2) user-defined nonzero values - useful for instances in which the pipeline needs to handle data dropouts in a different manner than actual zeros in a channel, or (3) the last good value - useful for channels containing constants that rarely change. Additionally, during real-time calibration, the pipeline uses a timer to determine when a dropout is occurring. Prior to this improvement, made during O3, the pipeline could only identify a dropout based on the timestamp of the first input data after the dropout ended. Long dropouts therefore caused the pipeline output to incur latency, causing the delayed strain data to be dropped in the data aggregation processes that follow calibration. This is no longer an issue, as the pipeline continuously fills in missing data during long dropouts.
@@ 4.1.3 Development of an Exact Solution for Time Dependence in the Calibration Model
The methods currently used to compute time-dependent parameters in $C^{\rm(model)}$ and $A^{\rm(model)}$ rely on simplifying approximations, resulting in $\sim$1% errors in the estimates of time-dependent parameters. The out-of-phase addition of the actuation and sensing components of $h_{\rm det}^{\rm(model)}$ shown in (6) inflates these errors in the strain data at some frequencies. An “exact” solution, described in chapter 5 of [59] has been developed in the gstlal calibration pipeline but was not yet implemented in the production of strain data during O4 as review and validation are still ongoing. This method will likely be implemented in O5, eliminating a systematic error that is currently uncompensated in the production of strain data. These systematic errors are largely accounted for—though not explicitly modeled—when estimating the overall calibration error and uncertainty budget in O4. The implementation of the exact solution may also produce more consistent improvements to calibration accuracy with the application of the full time-dependent model.
@@ 4.1.4 Subtraction of Spectral Lines
Searches for GWs are negatively impacted by narrow-band noise sources in the strain data [60]. Some narrow-band noise sources in the form of persistent, sinusoidal signals, known as calibration lines, are intentionally added via the Pcal lasers and the actuation system, in part to measure the time-dependent parameters of the calibration models during observation. A method for subtracting spectral lines, including the calibration lines, in low latency was implemented in the gstlal calibration pipeline starting at the beginning of O3. The method relies on the use of witness channels, which are correlated to the signal to be subtracted but insensitive to GWs. This is intended to prevent the unintentional subtraction of GWs. The method is carried out by first demodulating strain $h(t)$ and the witness channel $w_{j}(t)$ at the frequency of the $i^{\rm th}$ line $f_{i}$ to find complex values $\tilde{h}(f_{i})$ and $\tilde{w}_{j}(f_{i})$ which encode the amplitude and phase of $h$ and $w_{j}$ at that frequency. A transfer function value is then computed at $f_{i}$ using a running median as
@@
During the review of calibrated strain data in O3 and O4, several issues and shortfalls of the low-latency line subtraction performed by the calibration pipeline were identified. These issues were addressed and fixed before the final release of the O3 and O4 strain data. One problem that was addressed during O4 is a narrow-band systematic error, often referred to as “over-subtraction,” introduced near the line frequencies. This error is caused by the use of a finite amount of data to inform the transfer function values $T_{j}$, leading to an imperfect ability to distinguish between signal that is present in the witness $w_{j}$ and signal that is not. The bandwidth of this error is inversely proportional to the length of the running median, which had been set to 128 seconds during O3. Increasing the length of the running median mitigates the error, but it also hinders the ability to subtract spectral lines when the $T_{j}$ are evolving on shorter timescales. During O4, a method was implemented that allows the subtraction algorithm to dynamically change the length of the running median based on whether significant secular change is detected in the $T_{j}$ before taking the median. The length of the running median is now adjusted between a minimum of 128 seconds and a maximum of 4096 seconds. Figure 2 demonstrates the subtraction of the 17.1 Hz Pcal line using the 128 second running median, the 4096 second running median, and the adaptive switching method.
Numerous types of occasional line subtraction failures have been caused by data dropouts and by insufficient communication of detector state information to the subtraction algorithm. Many of these have been addressed by more sophisticated handling of data dropouts and by more thoroughly utilizing all available state information. The O4 implementation of line subtraction uses multiple metrics to determine when the input data is suitable for measurement of the $T_{j}$ and when the calibration lines are turned on or off. With the addition of the Pcal lines used for calibration monitoring in O4 (discussed below), the line subtraction performed in the low-latency calibration procedure has become increasingly essential to the production of high-quality strain data for downstream analyses. The improvements to the quality and robustness of the line subtraction algorithm have been essential to this achievement.